OpenStudy (anonymous):

Simplify the expression. –4.7 + 5.5

OpenStudy (anonymous):

you can subtract that separately: 5.5 - 4.7 = 5.5 - 4.0 - 0.7

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

.8

OpenStudy (anonymous):

4.7 = 4.0 + 0.7 -4.7 = -4.0 - 0.7

OpenStudy (anonymous):

do it step by step

OpenStudy (anonymous):

5.5 - 4.0 - 0.7

OpenStudy (anonymous):

the answer is .8

OpenStudy (anonymous):

thanks u guys

OpenStudy (anonymous):

.7 is incorrect.

OpenStudy (anonymous):

.7 is not the final answer.